1.9 10 Year Revenues Manufacturing Expenses (other than depreciation) Marketing Expenses Depreciation EBIT Taxes at 34% Unlevered Net Income Depreciation Additions to Net Working Capital Capital Expenditures Continuation Value Free Cash Flow 105.0 34.1 105.0 34.1 - 15.2 46.3 - 15.742- 15.742 30.558 15.2 - 15.2 46.3 30.558 15.2 - 152.0 + 11.6 53.158 - 152.0 41.558Bauer Industries is an automobile manufacturer. Management is currently evaluating a proposal to build a plant that will manufacture lightweight trucks. Bauer plans to use a cost of capital of 12.3 % to evaluate this project. Based on extensive​ research, it has prepared the following incremental free cash flow projections​ (in millions of​ dollars): a. For this​ base-case scenario, what is the NPV of the plant to manufacture lightweight​ trucks? b. Based on input from the marketing​ department, Bauer is uncertain about its revenue forecast. In​ particular, management would like to examine the sensitivity of the NPV to the revenue assumptions. What is the NPV of this project if revenues are 10 % higher than​ forecast? What is the NPV if revenues are 10 % lower than​ forecast? c. Rather than assuming that cash flows for this project are​ constant, management would like to explore the sensitivity of its analysis to possible growth in revenues and operating expenses.​ Specifically, management would like to assume that​ revenues, manufacturing​ expenses, and marketing expenses are as given in the table for year 1 and grow by 2 % per year every year starting in year 2. Management also plans to assume that the initial capital expenditures​ (and therefore​ depreciation), additions to working​ capital, and continuation value remain as initially specified in the table. What is the NPV of this project under these alternative​ assumptions? How does the NPV change if the revenues and operating expenses grow by 6 % per year rather than by 2 %​? d. To examine the sensitivity of this project to the discount​ rate, management would like to compute the NPV for different discount rates. Create a​ graph, with the discount rate on the x​-axis and the NPV on the y​-axis, for discount rates ranging from 5 % to 30 %. For what ranges of discount rates does the project have a positive​ NPV?
Answer #1

Since, there are multiple parts to the question and Part b and Part c have multiple subparts, I have answered the first three parts (a,b and c) completely..

______

Part a)

The NPV of the plant to manufacture lightweight​ trucks is calculated as below:

NPV = Capital Expenditures + Free Cash Flow Year 1/(1+Cost of Capital)^1 + Free Cash Flow Year 2/(1+Cost of Capital)^2 + Free Cash Flow Year 3/(1+Cost of Capital)^3 + Free Cash Flow Year 4/(1+Cost of Capital)^4 + Free Cash Flow Year 5/(1+Cost of Capital)^5 + Free Cash Flow Year 6/(1+Cost of Capital)^6 + Free Cash Flow Year 7/(1+Cost of Capital)^7 + Free Cash Flow Year 8/(1+Cost of Capital)^8 + Free Cash Flow Year 9/(1+Cost of Capital)^9 + Free Cash Flow Year 10/(1+Cost of Capital)^10

Using the values provided in the question in the above formula, we get,

NPV = -152 + 41.558/(1+12.3%)^1 + 41.558/(1+12.3%)^2 + 41.558/(1+12.3%)^3 + 41.558/(1+12.3%)^4 + 41.558/(1+12.3%)^5 + 41.558/(1+12.3%)^6 + 41.558/(1+12.3%)^7 + 41.558/(1+12.3%)^8 + 41.558/(1+12.3%)^9 + 53.158/(1+12.3%)^10 = $83.59 or $83.6 million

______

Part b)

When revenues are 10% higher than forecast

We will have to calculate revised free cash flow as below:

Year 0 1-9 10
Revenues (105 + 10%*105) 115.5 115.5
Manufacturing Expenses (Other than Depreciation) -34.1 -34.1
Marketing Expenses -9.4 -9.4
Depreciation -15.2 -15.2
EBIT 56.8 56.8
Taxes at 34% -19.312 -19.312
Unlevered Net Income 37.488 37.488
Depreciation 15.2 15.2
Additions to Net Working Capital -4.2 -4.2
Capital Expenditures -152
Continuation Value 11.6
Free Cash Flow -152 48.488 60.088

Now, we can calculate NPV as below:

NPV = -152 + 48.488/(1+12.3%)^1 + 48.488/(1+12.3%)^2 + 48.488/(1+12.3%)^3 + 48.488/(1+12.3%)^4 + 48.488/(1+12.3%)^5 + 48.488/(1+12.3%)^6 + 48.488/(1+12.3%)^7 + 48.488/(1+12.3%)^8 + 48.488/(1+12.3%)^9 + 60.088/(1+12.3%)^10 = $122.27 or $122.3 million

______

When revenues are 10% lower than forecast

We will have to calculate revised free cash flow as below:

Year 0 1-9 Years 10
Revenues (105 - 10%*105) 94.5 94.5
Manufacturing Expenses (Other than Depreciation) -34.1 -34.1
Marketing Expenses -9.4 -9.4
Depreciation -15.2 -15.2
EBIT 35.8 35.8
Taxes at 34% -12.172 -12.172
Unlevered Net Income 23.628 23.628
Depreciation 15.2 15.2
Additions to Net Working Capital -4.2 -4.2
Capital Expenditures -152
Continuation Value 11.6
Free Cash Flow -152 34.628 46.228

Now, we can calculate NPV as below:

NPV = -152 + 34.628/(1+12.3%)^1 + 34.628/(1+12.3%)^2 + 34.628/(1+12.3%)^3 + 34.628/(1+12.3%)^4 + 34.628/(1+12.3%)^5 + 34.628/(1+12.3%)^6 + 34.628/(1+12.3%)^7 + 34.628/(1+12.3%)^8 + 34.628/(1+12.3%)^9 + 46.228/(1+12.3%)^10 = $44.91 or $44.9 million

______

Part c)

When revenues and expenses grow by 2%

We will have to calculate revised free cash flow for each year as follows:

Year 0 1 2 3 4 5 6 7 8 9 10
Revenues 105.0 107.1 109.2 111.4 113.7 115.9 118.2 120.6 123.0 125.5
Manufacturing Expenses (Other than Depreciation) -34.1 -34.8 -35.5 -36.2 -36.9 -37.6 -38.4 -39.2 -40.0 -40.8
Marketing Expenses -9.4 -9.6 -9.8 -10.0 -10.2 -10.4 -10.6 -10.8 -11.0 -11.2
Depreciation -15.2 -15.2 -15.2 -15.2 -15.2 -15.2 -15.2 -15.2 -15.2 -15.2
EBIT 46.3 47.5 48.8 50.1 51.4 52.7 54.1 55.4 56.9 58.3
Taxes at 34% -15.742 -16.160 -16.587 -17.022 -17.466 -17.918 -18.380 -18.851 -19.331 -19.821
Unlevered Net Income 30.558 31.370 32.198 33.042 33.904 34.783 35.679 36.593 37.526 38.477
Depreciation 15.2 15.2 15.2 15.2 15.2 15.2 15.2 15.2 15.2 15.2
Additions to Net Working Capital -4.2 -4.2 -4.2 -4.2 -4.2 -4.2 -4.2 -4.2 -4.2 -4.2
Capital Expenditures -152
Continuation Value 11.6
Free Cash Flow -152 41.558 42.370 43.198 44.042 44.904 45.783 46.679 47.593 48.526 61.077

Now, we can calculate NPV as below:

NPV = -152 + 41.558/(1+12.3%)^1 + 42.370/(1+12.3%)^2 + 43.198/(1+12.3%)^3 + 44.042/(1+12.3%)^4 + 44.904/(1+12.3%)^5 + 45.783/(1+12.3%)^6 + 46.679/(1+12.3%)^7 + 47.593/(1+12.3%)^8 + 48.526/(1+12.3%)^9 + 61.077/(1+12.3%)^10 = $100.53 or $100.5 million

______

When revenues and expenses grow by 5%

We will have to calculate revised free cash flow for each year as follows:

Year 0 1 2 3 4 5 6 7 8 9 10
Revenues 105.0 110.3 115.8 121.6 127.6 134.0 140.7 147.7 155.1 162.9
Manufacturing Expenses (Other than Depreciation) -34.1 -35.8 -37.6 -39.5 -41.4 -43.5 -45.7 -48.0 -50.4 -52.9
Marketing Expenses -9.4 -9.9 -10.4 -10.9 -11.4 -12.0 -12.6 -13.2 -13.9 -14.6
Depreciation -15.2 -15.2 -15.2 -15.2 -15.2 -15.2 -15.2 -15.2 -15.2 -15.2
EBIT 46.3 49.4 52.6 56.0 59.6 63.3 67.2 71.3 75.7 80.2
Taxes at 34% -15.742 -16.788 -17.885 -19.038 -20.248 -21.519 -22.853 -24.254 -25.726 -27.270
Unlevered Net Income 30.558 32.588 34.718 36.956 39.305 41.772 44.362 47.082 49.938 52.936
Depreciation 15.2 15.2 15.2 15.2 15.2 15.2 15.2 15.2 15.2 15.2
Additions to Net Working Capital -4.2 -4.2 -4.2 -4.2 -4.2 -4.2 -4.2 -4.2 -4.2 -4.2
Capital Expenditures -152
Continuation Value 11.6
Free Cash Flow -152 41.558 43.588 45.718 47.956 50.305 52.772 55.362 58.082 60.938 75.536

Now, we can calculate NPV as below:

NPV = -152 + 41.558/(1+12.3%)^1 + 43.588/(1+12.3%)^2 + 45.718/(1+12.3%)^3 + 47.956/(1+12.3%)^4 + 50.305/(1+12.3%)^5 + 52.772/(1+12.3%)^6 + 55.362/(1+12.3%)^7 + 58.082/(1+12.3%)^8 + 60.938/(1+12.3%)^9 + 75.536/(1+12.3%)^10 = $129.15 or $129.1 million